OpenStudy (madinaxxx):

f(x) = 4x- 7; g(x) = x + 3 What is the value of (f * g)(4)? (Points : 4) 4 9 12 21

jimthompson5910 (jim_thompson5910):

g(x) = x + 3 g(4) = 4 + 3 ... replace every x with 4 g(4) = 7

jimthompson5910 (jim_thompson5910):

\[\Large (f \circ g)(x) = f(g(x))\]

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = 4x-7\] \[\Large f(g(x)) = 4(g(x))-7\] \[\Large f(g(4)) = 4(g(4))-7\] \[\Large f(g(4)) = 4(7)-7\] I'll let you finish up

OpenStudy (madinaxxx):

21?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (madinaxxx):

thanks

jimthompson5910 (jim_thompson5910):

np